Reduce the following fraction into partial fractions: 3x^ 2 + 7x + 28/x(x^ 2 + x + 7)


1
Expert's answer
2021-06-21T11:17:59-0400
"\dfrac{3x^2+7x+28}{x(x^2+x+7)}=\dfrac{A}{x}+\dfrac{B}{x^2+x+7}+\dfrac{Cx}{x^2+x+7}"

"=\dfrac{Ax^2+Ax+7A+Bx+Cx^2}{x(x^2+x+7)}"

"x^2: A+C=3"


"x^1: A+B=7"


"x^0: 7A=28"


"A=4"

"B=3"

"C=-1"



"\dfrac{3x^2+7x+28}{x(x^2+x+7)}=\dfrac{4}{x}+\dfrac{3}{x^2+x+7}-\dfrac{x}{x^2+x+7}"